OpenStudy (anonymous):

Solve for x: 2 over quantity x plus 2 plus 1 over 5 equals 6 over quantity x plus 5

OpenStudy (anonymous):

2/x + 2 + 1/5 = 6/x + 5 ?

OpenStudy (anonymous):

Si

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ 11 }{ 5 } - 5 = \frac{ 6 }{ x } - \frac{ 2 }{ x }\]\[\frac{ -14 }{ 5 } = \frac{ 4 }{ x }\] \[\frac{ x }{ 5 } = \frac{ 4 }{-14}\] \[x = -\frac{ 4 }{ 14 } \times 5\] \[x = -\frac{ 10 }{ 7 }\]

OpenStudy (anonymous):

my approach is all the x terms in the left hand side and others on right hand side

OpenStudy (anonymous):

if you dont understand some parts of the work out just ask me ill help you :)

OpenStudy (anonymous):

THX